OpenStudy (anonymous):

Bob pays $20 to have the Internet installed and $15 a month for service. How much will Bob pay for 1 year of Internet use? A. $180 B. $200 C. $240 D. $420

OpenStudy (anonymous):

20+15m=? 20+15(12)=? 20+180=? 20+180=200

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

You're welcome.